OpenStudy (anonymous):

how do I solve the problem: Za/2 corresponds with a 85% CL

jimthompson5910 (jim_thompson5910):

You need to find the value of k such that P(Z < -k) = 0.075 This is because 85% (or 0.85) of the distribution is between -k and k. Where k is some positive number. The remaining amount of the distribution, which is 100 - 85 = 15%, is divided equally between the tails. So each tail gets 15/2 = 7.5% So that's how I got 0.075

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

Yes, it does! Now why couldn't my book explain it in layman's terms like that?! =) thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Ok, so my HW says that's wrong.. looking for a level of confidence = (1-a) *100% .85 = (1-a)*1

jimthompson5910 (jim_thompson5910):

what did you get for that value of k

jimthompson5910 (jim_thompson5910):

and if possible, post a screenshot of the entire problem please

OpenStudy (anonymous):

I don't know how to do a screen shot..nor solve for K.

OpenStudy (anonymous):

or how k ended up in the equation..

jimthompson5910 (jim_thompson5910):

you hit print screen open up MS paint paste it in save it, send it

OpenStudy (anonymous):

'm using a mac

jimthompson5910 (jim_thompson5910):

well k is an added variable

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

mac OS X right?

OpenStudy (anonymous):

affirmative.

jimthompson5910 (jim_thompson5910):

you hit Command-Shift-3 that will save a PNG file to the desktop

jimthompson5910 (jim_thompson5910):

the PNG file will be what you are currently looking at

jimthompson5910 (jim_thompson5910):

ok you want to find the z score such that the area to the left of the z-score is 0.075 so if k is that z-score (adding this variable in) then P(Z < k) = 0.075

jimthompson5910 (jim_thompson5910):

this part is what was found above

jimthompson5910 (jim_thompson5910):

or if you have a TI calculator, you can use the "invNorm" function and type in invNorm(0.075) to get -1.44 roughly

jimthompson5910 (jim_thompson5910):

so -k = -1.44 k = 1.44

jimthompson5910 (jim_thompson5910):

Which means that the z critical value for the 85% confidence interval is roughly 1.44

OpenStudy (anonymous):

where is the "invNorm" function roughly located on the TI calculator?

OpenStudy (anonymous):

brb

jimthompson5910 (jim_thompson5910):

hit 2nd key, then the VARS key

jimthompson5910 (jim_thompson5910):

it's option #3

OpenStudy (anonymous):

got it.

jimthompson5910 (jim_thompson5910):

ok so you were able to get -1.44 or -1.439531471... right?

OpenStudy (anonymous):

yes, which you divide by 1 to get the positive #?

OpenStudy (anonymous):

or rather the -1

jimthompson5910 (jim_thompson5910):

the positive version of that number is 1.439531471 so that's your z critical for the 85% CL

OpenStudy (anonymous):

ha. that made since but I input the answer wrong again..and it gave me the correct answer as to being .15..did I need to move a decimal?

jimthompson5910 (jim_thompson5910):

it says the answer is 0.15 ?

OpenStudy (anonymous):

yes. I'll screen shot it.

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

oooh it wants the alpha value

jimthompson5910 (jim_thompson5910):

if CL = 0.85 then alpha = 1 - CL alpha = 1 - 0.85 alpha = 0.15

jimthompson5910 (jim_thompson5910):

but that's odd since you wanted the critical value first, but oh well

OpenStudy (anonymous):

oooh.

jimthompson5910 (jim_thompson5910):

oh you're still doing the example, I gotcha

jimthompson5910 (jim_thompson5910):

sorry wasn't sure where you were at lol

OpenStudy (anonymous):

that's ok. I'm not sure of that either. =)

jimthompson5910 (jim_thompson5910):

but anyway on the main problem for #2, you find the alpha value (in this case 0.15) and you use that to determine z(a/2) which is the critical value for the confidence interval

jimthompson5910 (jim_thompson5910):

I'm sure that this example will have you using a table, but a calculator is quicker, easier, and more accurate

OpenStudy (anonymous):

to my knowledge, no tables were provided..let me double check my book.

jimthompson5910 (jim_thompson5910):

alright, tables might be provided in the front or back of the book they also might be given out as extra stuff to download (or maybe given out as links)

jimthompson5910 (jim_thompson5910):

I know I have a few tables in PDF form (just the tables themselves) on my computer when I took stats since the professor sent a few that way

OpenStudy (anonymous):

nope, I don't see anything.

jimthompson5910 (jim_thompson5910):

ok then maybe your teacher has posted the tables or has sent the tables in another form

jimthompson5910 (jim_thompson5910):

sometimes, when doing a problem like this, the table will be given alongside the problem (like you click a button off to the side of the problem and the table used will pop up)

OpenStudy (anonymous):

nope, no table provided. We had the answer but the question was breaking it down step by step..and I was jumping ahead. thank you for your assistance. I'm going to try another problem and see how it goes.

jimthompson5910 (jim_thompson5910):

alright, well anyway, if you break from the example in #2 and go back to the main problem #2 the final answer should be 1.43953147, make sure you round this to whatever place they want

OpenStudy (anonymous):

yes, I did..and it was correct - woohoo!

jimthompson5910 (jim_thompson5910):

alright great lol